rosemaryj1401:

Box A has a volume of 12 cubic meters. Box B is similar to Box A. To create Box B, Box A's dimensions were multiplied by five. What is the volume of Box B?

harliii:

the answer is 60 Well I think Since the dimensions of A were multiplied by 5 to create B then you will have to multiply the volume of Box A by 5 in order to find the volume of Box B Box B volume = 5 * box A volume Box B volume = 5 * 12 Box B volume = 60

Mercury:

old question but will answer anyway because a box has three dimensions, if each dimension is multiplied by 5 the volume is multiplied by 5*5*5 not just 5